How many grams of Aluminum Bromide would be required to react with 0.10 moles of tin metal?


1
Expert's answer
2021-03-22T06:28:13-0400

Molar Mass of tin metals = 118 g/mol

Molar Mass of aluminum bromide = 266g/mol

0.1 mol of tin = 11.871g

= X × (1 mol Sn/11.871) × (2mol/1mol) = 0.1 mol

X = 0.1×5.9

= 0.59g of Aluminum bromide